base of a triangle is 7 feet more than 1/2 height, find height if area of triangle is 170 square feet
Answer by checkley77(12844)

B=H/2+7
A=BH/2
170=(H/2+7)*H/2
170=H^2/4+7H/2
170=(H^2+2*7H)/4
H^2+14H=4*170
H^2=14H=680
H^2+14H-580=0
H-20)(H+34)=0
H-20=0
H=20 FT.
PROOF:
170=(20/2+7)*20/2
170=(10*7)*10
170=17*10
170=170